OpenStudy (anonymous):

The sum of an angle's complementary and supplementary is 124. Find the measure of the angle.

OpenStudy (anonymous):

I can't make up the equation. Please help

OpenStudy (anonymous):

i think 180- 124 = 56

OpenStudy (anonymous):

why would that be the equation?

OpenStudy (anonymous):

u want the measure of remaining angle right? then ans is 56

OpenStudy (anonymous):

lets call the angle a complementary of a = 90-a supplementary of a = 180-a their sum is 124: 90-a + 180-a = 124 solve for a

OpenStudy (anonymous):

that's what i thought! thanks!

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

thank you for correcting me